BigCo Inc. currently has a bond issue outstanding with 8 years to maturity, a face value of $1,000 and a coupon rate of 7.00% (paid annually). The bond currently sells for $1,062. BigCo’s common stock has a beta of 0.85, and currently sells for $56/share. Treasury bills yield 2.50% and the expected rate of return on the market is 9%. BigCo’s marginal tax rate is 21%. BigCo’s total bond debt has a face value of $8,000,000 and there are 400,000 shares of stock outstanding. You must show calculator keystrokes or equations to receive any credit. No points for the final answer only.

Estimate BigCo’s nominal (pre-tax) cost of debt:

Estimate BigCo’s cost of equity:

What is BigCo’s WACC?

If BigCo has a new project that costs $2.5 million today and will return a one-time cash flow of $4.2 million in one year. If this project is similarly risky as the existing firm’s assets, what is the project’s NPV?
Answer #1

1.
N=8
PMT=7%*1000
PV=-1062
FV=1000
CPT I/Y=6.00%

2.
=2.50%+0.85*(9%-2.50%)
=8.03%

3.
=(8000000*1062/1000*6%*(1-21%)+400000*56*8.03%)/(8000000*1062/1000+400000*56)
=7.13%

4.
=-2.5+4.2/(1+7.13%)
=1.42047 million